40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jun 20 00:59:02 2011, kiwamu, Summary, General, Weekly report 
    Reply  Mon Jun 27 10:04:29 2011, kiwamu, Summary, General, Weekly report 
       Reply  Mon Jul 4 14:27:35 2011, kiwamu, Summary, General, Weekly report 
          Reply  Mon Jul 25 11:02:19 2011, kiwamu, Summary, General, Weekly report 
             Reply  Mon Jul 25 11:04:22 2011, kiwamu, Summary, General, Weekly report 
                Reply  Mon Jul 25 11:06:38 2011, kiwamu, Summary, General, Weekly report 
                   Reply  Mon Aug 1 22:30:59 2011, kiwamu, Summary, General, Weekly report 
Message ID: 5085     Entry time: Mon Aug 1 22:30:59 2011     In reply to: 5028
Author: kiwamu 
Type: Summary 
Category: General 
Subject: Weekly report 

Summary of the week ending July 31st.  Number of elog entries = 53


- SUS
 + ETMY-LR sensor looked strange. Something wrong.
 + Responses from the DC alignment bias to the shadow sensors and the oplevs were checked.
   --> ETMY shows the response with the opposite sign. Wired.
 + ETMY shadow sensors were examined in terms of the spectra.
   --> WF, AA and ADC noise looks reasonably low and not high enough to explain the low frequency noise.
 + Adjusted all the OSEM gains

- LSC
 + MICH noise budget is ongoing. WF filter needs to be greater than 21 dB to have dark noise of the PD greater than ADC noise
 + The arms became lockable

- CDS
 + modified and re-ran activateDQ.py.
 + c1iscex crashed for unknown reasons and we physically rebooted it.

- ALS (Fiber experiments)
 + PMC trans is sampled for the fiber beat-note measurement on the PSL table.
 + The beat-note signal between PSL and X end laser were obtained.
 + Some optics in the ETMY table were rearragend to have the Y green light aligned.

- ACS / ASS
 + incident beam axis has changed a lot.
 + X arm and Y arm ASS were reactivated.
  ---> The sign of some of the control gains had been wrong.
 + The incident beam axis and X/Y arm were re-aligned

- IOO/MCWFS
 + Some medm screens fixed.
 + Adjustment of the demodulation phase on each quadrant on WFS1 and WFS2 are done.
 + The sensing matrix (from optics to WFS sensors) were measured.

- OAF
 + c1pem was modified
 + plugged a seismometers to ADC through an AA board.
   --> channels are coming to the digital land
 
- Preparation for the invac work
 + 7 pieces of beam traps are available
 + Tolerance of the arm length is estimated to be +/- 2 cm.


- PSL/RefCav
 + ABSL is injected into the reference cavity. some flashing happened but no locking.
 + eddited the psl.db file to set EGUF and EGUL
 + turned RefCav heater and servo back on

ELOG V3.1.3-